# Constraining Schwarzschild Models with Orbit Classifications

### v1

@@ 5.2 Orbit Classifications from Grid-searching
When no target orbit classification is available, it is possible to construct a grid search of Schwarzschild models to determine an appropriate classification from the orbit data. The search used here is to construct classifications by randomly uniformly selecting values from supplied minimum to maximum class ranges, and then to score those classifications using unregularized Schwarschild modeling. A number of classifications closest in (Euclidean) distance to the classification with the minimum score are used to set revised minimum to maximum class ranges, and the classification construction and scoring processes are repeated. For the simulated galaxies, two iterations after the initial scoring are sufficient to have converged to an orbit classification for a galaxy.
@@
Row ‘grid search’ indicates the orbit classifications determined using the grid search approach. Row ‘without reg’ repeats the orbit classifications from Tables 3 and 6. For the simulated galaxies, the orbit classifications differ by at most $0.01$. Some of the differences for the ATLAS3D galaxies are $>0.01$, perhaps indicating that a 3rd search iteration was needed due to the wider cold and counter-rotating ranges.
@@
5.3 Initial Conditions - Rotation Extremes
In Section 5.1, all the models have initial conditions where there is a sense of bulk rotation aligned with the mean velocity data. In Table 6, a set of results for ATLAS3D galaxy models is shown where the models are lacking an overall sense of rotation. The actual ATLAS3D galaxies do have a sense of rotation - see Table 1.
Table 6: Results for ATLAS3D Galaxies - no overall sense of rotation
@@ 5.3 Initial Conditions - Rotation Extremes
In summary, where the kinematical data exhibit bulk rotation of a galaxy, full aligned rotation is unlikely to be satisfactory: some counter rotating orbits appear to be necessary to reproduce the observed data. No sense of rotation may be appropriate if no overall rotation or only very minor rotation is observed. This leaves modeling of many galaxies in a position where some overall sense of rotation should be applied to the initial conditions but just how much ought to be determined via experimentation.
5.4 Initial Conditions - An Alternative to 3I
An alternative scheme for initial conditions is the MDJV match density Jeans velocities scheme described in Long et al. (2021). The motivation for MDJV is to use the observed data in setting the initial conditions. A set of models of the simulated galaxies using MDJV is constructed and compared with the equivalent 3I baseline models in Section 5.1.
@@ 5.4 Initial Conditions - An Alternative to 3I
Comparing the MDJV results in Table 7 with the 3I results in Table 3, the mean model $\chi^{2}$ values for the MDJV models are all lower than their 3I counterparts indicating that overfitting has increased. The maximum mean $\chi^{2}$ value for MDJV is 0.29 while the maximum for 3I is 0.64. The number of active orbits is increased with the MDJV models. From the results achieved, there is no reason to doubt the validity of using MDJV initial conditions. However MDJV does produce orbit classifications that do not agree with those produced by the 3I scheme.
@@ 5.5 Removal of Orbit Classes
With the removal of cold orbits from the ATLAS3D models for galaxies [NGC 1248, NGC 3838, NGC 4452, NGC 4551], the mean model $\chi^{2}$ values are $[0.34,0.71,1.39,0.21]$. Observable reproduction for NGC 4452 is close to unacceptable (all observable mean $\chi^{2}$ are $>1$; there are no concerns for the other three galaxy models.
@@ 6 Discussion
As set out in the Introduction, Section 1, the science matter is concerned with how to gain knowledge of the internal stellar orbital structure of external galaxies in the spite of the inability to obtain 6D phase space data, and to observe stellar orbits. Astronomers have only blended observational data to rely on to try and gain knowledge of structures that can not be directly validated.
@@
The advent of large language models (LLMs), a type of generative AI (artificial intelligence) application, introduces another complication. Astronomers have to be very careful about caveating their results appropriately and not, for example, overstating their results. Astronomy articles form part of the input to LLMs, and new researchers already risk being misinformed by unwisely formulated science articles.
@@ 7 Conclusions
The objective set out in the Introduction, Section 1, has been met. The question posed was ‘could an orbit classification for a galaxy being investigated be included in Schwarzschild modeling as a constraint’ ? As has been demonstrated, the answer is ’yes’. Unfortunately, quite where that orbit classification might be obtained from is an open question: it does not appear to be from using Schwarzschild’s method given its usual implementation; it might be from cosmological simulations. Having the capability to include classification constraints is, however, another useful technique in the stellar dynamical toolbox.
@@ Acknowledgements.
The author thanks Dandan Xu, Ling Zhu and Shude Mao for various fruitful discussions as the research recorded here progressed. This work is partly supported by the National Key Basic Research and Development Program of China (No. 2018YFA0404501 to Shude Mao), and by the National Science Foundation of China (Grant No. 11821303, 11761131004 and 11761141012 to Shude Mao).

### v2

@@ 5.2 Orbit Classifications from Grid-searching
When no target orbit classification is available, it is possible to construct a grid search of Schwarzschild models to determine an appropriate classification from the orbit data. The search used here is to construct classifications by randomly uniformly selecting values from supplied minimum to maximum class ranges, and then to score those classifications using unregularized Schwarzschild modeling. A number of classifications closest in (Euclidean) distance to the classification with the minimum score are used to set revised minimum to maximum class ranges, and the classification construction and scoring processes are repeated. For the simulated galaxies, two iterations after the initial scoring are sufficient to have converged to an orbit classification for a galaxy.
@@
Row ‘grid search’ indicates the orbit classifications determined using the grid search approach. Row ‘without reg’ repeats the orbit classifications from Tables 3 and 6. For the simulated galaxies, the orbit classifications differ by at most $0.01$. Some of the differences for the ATLAS3D galaxies are $>0.01$, perhaps indicating that a 3rd search iteration is needed due to the wider cold and counter-rotating ranges.
@@
5.3 Initial Conditions—Rotation Extremes
In Section 5.1, all the models have initial conditions where there is a sense of bulk rotation aligned with the mean velocity data. In Table 6, a set of results for ATLAS3D galaxy models is shown where the models are lacking an overall sense of rotation. The actual ATLAS3D galaxies do have a sense of rotation—see Table 1.
Table 6: Results for ATLAS3D Galaxies—No Overall Sense of Rotation
@@ 5.3 Initial Conditions—Rotation Extremes
In summary, where the kinematical data exhibit bulk rotation of a galaxy, fully aligned rotation is unlikely to be satisfactory: some counter-rotating orbits appear to be necessary to reproduce the observed data. No sense of rotation may be appropriate if no overall rotation or only very minor rotation is observed. This leaves modeling of many galaxies in a position where some overall sense of rotation should be applied to the initial conditions but just how much ought to be determined via experimentation.
5.4 Initial Conditions—An Alternative to 3I
An alternative scheme for initial conditions is the MDJV, match density Jeans velocities, scheme described in Long et al. (2021). The motivation for MDJV is to use the observed data in setting the initial conditions. A set of models of the simulated galaxies using MDJV is constructed and compared with the equivalent 3I baseline models in Section 5.1.
@@ 5.4 Initial Conditions—An Alternative to 3I
Comparing the MDJV results in Table 7 with the 3I results in Table 3, the mean model $\chi^{2}$ values for the MDJV models are all lower than their 3I counterparts indicating that overfitting has increased. The maximum mean $\chi^{2}$ value for MDJV is 0.29 while the maximum for 3I is 0.64. The number of active orbits is increased with the MDJV models. From the results achieved, there is no reason to doubt the validity of using MDJV initial conditions. MDJV does produce orbit classifications that do not agree with those produced by the 3I scheme.
@@ 5.5 Removal of Orbit Classes
With the removal of cold orbits from the ATLAS3D models for galaxies [NGC 1248, NGC 3838, NGC 4452, NGC 4551], the mean model $\chi^{2}$ values are $[0.34,0.71,1.39,0.21]$. Observable reproduction for NGC 4452 is close to unacceptable (all observable mean $\chi^{2}$ are $>1$); there are no concerns for the other three galaxy models.
@@ 6 Discussion
As set out in the Introduction, Section 1, the science matter is concerned with how to gain knowledge of the internal stellar orbital structure of external galaxies in spite of the inability to obtain 6D phase space data, and to observe stellar orbits. Astronomers have only blended observational data to rely on to try and gain knowledge of structures that cannot be directly validated.
@@
The advent of large language models (LLMs), a type of generative artificial intelligence application, introduces another complication. Astronomers have to be very careful about caveating their results appropriately and not, for example, overstating their results. Astronomy articles form part of the input to LLMs, and new researchers already risk being misinformed by unwisely formulated science articles.
@@ 7 Conclusions
The objective set out in the Introduction, Section 1, has been met. The question posed was ‘could an orbit classification for a galaxy being investigated be included in Schwarzschild modeling as a constraint?’ As has been demonstrated, the answer is ’yes.’ Unfortunately, quite where that orbit classification might be obtained from is an open question: it does not appear to be from using Schwarzschild’s method given its usual implementation; it might be from cosmological simulations. Having the capability to include classification constraints is, however, another useful technique in the stellar dynamical toolbox.
@@ Acknowledgements.
The author thanks Dandan Xu, Ling Zhu and Shude Mao for various fruitful discussions as the research recorded here progressed. This work is partly supported by the National Key Basic Research and Development Program of China (No. 2018YFA0404501 to Shude Mao), and by the National Natural Science Foundation of China (NSFC, grant Nos. 11821303, 11761131004 and 11761141012 to Shude Mao).